OpenStudy (anonymous):

Verify the identity Sec^6x(TanxSecx)- Sec^4(TanxSecx)= Sec^5xTan^3x

OpenStudy (anonymous):

\[\sec ^{6}x \left( \tan x \sec x \right)-\sec ^{4}x \left( \tan x \sec x \right)=\tan x \sec x \sec ^{4}x \left( \sec ^{2}x-1 \right)\] \[=\tan x \sec ^{5}x \tan ^{2}x=R.H.S.\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw